OpenStudy (anonymous):

If Sally can paint a house in 4 hours, and John can paint the same house in 6 hours, how long will it take both of them to paint the house together?

OpenStudy (amistre64):

the trick to solving this type of problem is to find out how much work is done by each person in 1 hour. Since S can paint one whole house in 4 hours; how much of the house cah she paint in 1 hour? S paints 1/4 of the house in 1 hour. How long does it take J to paint the entire house? ^ hours. which means that in 1 hour J has 1/6 of the house painted. How we can add together S and J to see how much of the house they can paint together in 1 hour: 1/4 + 1/6 = amount painted by both in 1 hour. 10/24, which reduces to 5/12 of the house is painted in one hour. then its just a matter of counting out the hours till it is completed... 1 hour = 5/12 2 hours = 10/12 there is only 2/12 of the house left to paint so its gonna take less than an hour... 2/12 reduces to 1/6; we divide the hour up into more basic units...into minutes. 60 minutes equals 1 hour and we want 1/6the of those minutes.. 60/6 = 10 minutes They can paint the house in 2 hours and 10 minutes.....